OpenStudy (anonymous):

The relation between time and velocity is distance(s)=3t+2t^2. Find Acceleration.

OpenStudy (anonymous):

Differentiate...) Two times

OpenStudy (anonymous):

v = ds/dt = 3 + 4t a = dv/dt = 4 m/s^2

OpenStudy (anonymous):

ds/dt? wat is d? and how did it became 4t

OpenStudy (anonymous):

Please Respond

OpenStudy (anonymous):

what's ur problem?

OpenStudy (anonymous):

how to find acceleration whent his relayion is given.

OpenStudy (anonymous):

s=3t+2t^2 s---distance traveled

OpenStudy (anonymous):

how to find acceleration

OpenStudy (anonymous):

yes s=distance